OpenStudy (anonymous):

find the distance between (4,2) and (-4,-4)

jimthompson5910 (jim_thompson5910):

Are you familiar with the distance formula?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

what do you get when you use that formula

jimthompson5910 (jim_thompson5910):

If you need help getting started, then you would do something like this... d = sqrt((x2-x1)^2+(y2-y1)^2) d = sqrt((-4-4)^2+(-4-2)^2) d = sqrt((-8)^2+(-6)^2)